OpenStudy (anonymous):

Jean says that you can take the square root of each side of an equation. Therefore, she decides that (x - 2)2 + (y + 3)2 = 36 and (x - 2) + (y + 3) = 6 are equuivalent equations. Marco says that the equations are not equivalent. Who is right? Explain your answer.

OpenStudy (across):

This is not true. :P

OpenStudy (hoblos):

not equivalent (x - 2) + (y + 3) = 6 if you square both sides you get (x - 2)^2 + (y + 3)^2 +2(x-2)(y+3) = 36

OpenStudy (across):

Take, for example,\[2^2+3^2=4+9=13\]\[2+3=5\]and\[\sqrt{13}\neq5\]